# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ Abstract
This article introduces a leave-one-out regression adjustment estimator (LOORA) for estimating average treatment effects in randomized controlled trials. The method removes the finite-sample bias of conventional regression adjustment and provides exact variance expressions for LOORA versions of the Horvitz-Thompson and difference-in-means estimators under simple and complete random assignment. Ridge regularization limits the influence of high-leverage observations, improving stability and precision in small samples. In large samples, LOORA attains the asymptotic efficiency of regression-adjusted estimator as in [lin2013agnostic], while remaining exactly unbiased. To construct confidence intervals, we rely on asymptotic variance estimates that treat the estimator as a two-step procedure, accounting for both the regression adjustment and the random assignment stages. Two within-subject experimental applications that provide realistic joint distributions of potential outcomes as ground truth show that LOORA eliminates substantial biases and achieves close-to-nominal confidence interval coverage.
@@ 1 Introduction
Establishing causal relations is a central goal in many sciences, as well as economics and the social sciences. With observational data, selection bias often obscures the true effect of interventions. Randomized controlled trials (RCTs) overcome this problem by assigning treatments randomly, eliminating systematic differences between treatment and control groups, and enabling unbiased comparisons of outcomes. RCTs now stand at the core of modern empirical research in causal inference.
@@
Most theoretical guarantees for regression-adjusted estimators rely on asymptotic arguments. However, early-stage clinical trials, marketing geo-experiments, and studies of firm- or country-level policies frequently operate in small sample regimes, where asymptotic results do not apply. In these cases, standard regression-adjusted estimators become biased, with undesirable consequences for treatment effect estimation and policy design. Moreover, high-leverage observations (that is, observations whose covariates strongly influence the regression fit) further degrade the performance of regression adjustment in small samples. This problem is well documented in influential work by freedman2008regression and young2019channeling.
@@
To address these challenges, in this paper we develop a regression adjustment framework that is unbiased in finite samples and robust to influential observations. A key component of our analysis is a leave-one-out regression adjustment (LOORA) procedure, which removes the bias of classical regression adjustment. Within this framework, we develop two estimators of the average treatment effect (ATE): LOORA-HT, a Horvitz–Thompson estimator for simple random assignment, and LOORA-DM, a difference-in-means estimator for complete random assignment. LOORA-HT and LOORA-DM are easy to implement and deliver a performance that previously required complex assignment mechanisms (see harshaw2022design).
@@
LOORA-HT and LOORA-DM simultaneously address three core problems in regression-adjusted estimation of treatment effects. First, they answer the critique of freedman_2008_b by providing unbiased, randomization-based estimators. Second, they reach the asymptotic efficiency of lin2013agnostic. Third, by stabilizing estimation in the presence of influential observations, they resolve the issues identified by young2019channeling.
@@
1.1 Related Work
The practice of regression adjustment in randomized experiments dates at least to the use of analysis of covariance (ANCOVA) in classical experimental design, the effects of pre-treatment covariates are removed through linear adjustment in order to improve precision. [^1] See, for example, discussions in fisher1971design. Despite its pervasiveness, the formal large-sample properties of regression adjustment after randomization have been repeatedly questioned. In two influential critiques, freedman2008regression; freedman_2008_b argue that ordinary least squares (OLS) adjustment can introduce bias (in both point estimates and standard errors) when the linear working model is misspecified, and can perform worse than the unadjusted difference-in-means estimator in finite samples.
In response, lin2013agnostic shows that, under complete random assignment, an OLS regression of outcomes on treatment, covariates, and all treatment-by-covariate interactions delivers an estimator that is consistent for the average treatment effect and asymptotically is at least as efficient as the unadjusted difference-in-means estimator, even when the linear model is misspecified. These results clarify the large sample properties of regression adjustment after randomization. However, two gaps remain in the literature: (i) lin2013agnostic does not provide a design-based finite-sample unbiasedness guarantee for regression adjustment, and (ii) the interacted specification in lin2013agnostic can be unstable in moderate samples because it increases the number of regressors, adding with additional interaction terms and potentially inflating leverage scores and sensitivity to influential units.
Our paper contributes to this literature by constructing estimators—LOORA-HT under simple random assignment and LOORA-DM under complete random assignment—that (i) are exactly unbiased for the finite-population average treatment effect, (ii) admit closed-form variance expressions in finite samples, and (iii) are asymptotically efficient, attaining the variance bound of lin2013agnostic. A key technical component is leave-one-out regression adjustment, in which each unit’s outcome is adjusted using coefficients estimated from a regression that excludes that unit. This leave-one-out procedure removes the finite-sample bias identified by freedman2008regression; freedman_2008_b. In addition, LOORA-HT and LOORA-DM use of ridge regression to control the impact of high-leverage observations. For the difference-in-means setting, we show that LOORA-DM admits a representation that is equivalent to a leave-two-out construction in the sense of spiess2025optimal, thereby satisfying necessary conditions for unbiasedness under complete random assignment. In this way, our estimators provide constructive finite-sample analogues of the asymptotic guarantees in lin2013agnostic, while remaining valid without requiring parametric outcome models.
@@ 1.1 Related Work
A related line of work examines precision gains in randomized studies through experimental design rather than regression adjustment. harshaw2022design propose a Gram–Schmidt walk design, which chooses treatment assignments in a way that approximately balances covariates and, in turn, controls the variance of the Horvitz–Thompson estimator through ridge-regularization. Their goal is to optimize the randomization design to deliver low-variance Horvitz–Thompson estimators in finite samples.
In contrast, we keep the assignment mechanism fixed (simple or complete random assignment) and instead adjust outcomes through leave-one-out regressions. This distinction matters in practice, because our methods can be applied to commonly used experimental designs without the need to change the assignment rule.
ghadiri2023finite introduce a cross-fitted regression-adjusted Horvitz–Thompson estimator under simple random assignment. Their estimator splits the sample into two folds, estimates regression coefficients on one fold, and applies those coefficients to adjust outcomes in the other fold, yielding an exactly unbiased estimator with an explicit nonasymptotic variance bound. However, the bound in ghadiri2023finite is looser than the lin2013agnostic asymptotic efficiency bound (by roughly a constant factor) and they do not develop confidence intervals.
Our leave-one-out construction both tightens the variance bound and delivers an estimator whose asymptotic variance matches lin2013agnostic. We also provide a simple two-step regression view that yields a Huber–White style variance estimator and confidence intervals that nearly match the nominal coverage values on two applications.
lei2021regression propose a bias-correction procedure for regression adjustment when the number of covariates grows with the sample size. Their method removes higher-order bias terms but does not yield exact unbiasedness in finite samples. chang2024exact develop an exact bias-correction approach that eliminates bias entirely, although they do not provide a finite-population variance bound for the resulting estimator. armstrong2021finite study the construction of optimal confidence intervals in finite samples under the assumption that regression errors are normally distributed with known variances. Finally, kline2020leave analyze leave-one-out methods for variance estimation in linear models, providing asymptotic results but not nonasymptotic guarantees.
@@
A large concurrent literature studies regression adjustment and inference in more structured designs: stratified or blocked experiments, matched pairs, cluster randomization, or covariate-adaptive randomization. For example, bai2024primer survey modern design-based analysis of randomized experiments, emphasizing the role of stratification, regression adjustment and cluster randomized experiments. bai2025new; bai2025inference develop design-based variance estimators and inference procedures for finely stratified and matched-pair experiments, including settings with imperfect compliance. cytrynbaum2024covariate characterizes, for general stratified randomization schemes, the asymptotically optimal linear covariate adjustment and shows that the classical interacted regression of lin2013agnostic can be inefficient away from complete random assignment. We also note that stratification becomes impractical in the presence of high-dimensional covariates, whereas our regression-adjustment approach remains applicable in such settings.
@@ 1.2 Notation
We denote matrices and vectors by bold uppercase and lowercase letters, respectively. The $i^{\text{th}}$ entry of a vector $\bm{u}$} is denoted by $u_{i}$. The transposed $i^{\text{th}}$ row of a matrix $\bm{X}$} is denoted by ${\bm{x}}_{i}$, and its $(i,j)^{\text{th}}$ entry by $x_{ij}$. For a constant $c$ and a vector $\bm{u}$}, expressions such as $c+{\bm{u}}$ and ${\bm{u}}^{-1}$ are interpreted entrywise. When $\bm{u}$} and $\bm{v}$} are vectors of equal dimension, ${\bm{u}}{\bm{v}}$ and ${\bm{u}}/{\bm{v}}$ denote entrywise (Hadamard) product and division, respectively. A vector’s associated diagonal matrix appears in uppercase; for instance, $\boldsymbol{T}$ denotes the diagonal matrix $\bm{t}$} as its main diagonal. For any vector ${\bm{v}}\in{\mathbb{R}}^{n}$, the notation ${\bm{v}}_{-i}$ refers to the vector in ${\mathbb{R}}^{n-1}$ obtained by removing the $i^{\text{th}}$ entry of $\bm{v}$}. Likewise, ${\bm{X}}_{-i}$ denotes the matrix obtained by deleting the $i^{\text{th}}$ row of $\bm{X}$}.
Norms and asymptotic notation.
For a vector ${\bm{u}}\in{\mathbb{R}}^{n}$, $\left\|{\bm{u}}\right\|_{2}=\sqrt{{\bm{u}}^{\top}{\bm{u}}}$ denotes the Euclidean norm, and $\left\|{\bm{u}}\right\|_{\infty}=\max_{i\in[n]}|u_{i}|$ denotes the $\ell_{\infty}$ norm. The operator ${\rm tr}(\cdot)$ denotes the trace of a square matrix. For a matrix ${\bm{X}}\in{\mathbb{R}}^{n\times k}$, $\left\|{\bm{X}}\right\|_{F}=\sqrt{{\rm tr}({\bm{X}}^{\top}{\bm{X}})}$ denotes the Frobenius norm, and $\left\|{\bm{X}}\right\|_{2,\infty}=\max_{i\in[n]}\left\|{\bm{x}}_{i}\right\|_{2}$ denotes its $(2,\infty)$-operator norm. Let $\mathbb{N}$ denote the set of natural numbers. For any two functions $f,g\colon\mathbb{N}\to{\mathbb{R}}$, we write $f(n)=O(g(n))$ if there exist constants $C>0$ and $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$, $\lvert f(n)\rvert\leq C\lvert g(n)\rvert$.
Projection matrices.
Consider a full-rank matrix ${\bm{X}}\in{\mathbb{R}}^{n\times k}$ with $n\geq k$. We denote the projection matrix by $\boldsymbol{H}={\bm{X}}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{X}}^{\top}$, and its diagonal entries by $h_{ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{x}}_{i}$, which represent the leverage scores of rows $i\in[n]$. The identity matrix is denoted by $\bm{I}$}. For $\lambda\geq 0$, the ridge projection matrix is $\boldsymbol{H}_{\lambda}={\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}$, whose diagonal entries $h_{\lambda ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}$ are the ridge leverage scores.
@@ 2 Finite Population Framework
We operate within the Neyman–Rubin potential outcomes framework (neyman1923application; rubin1974estimating) for a randomized controlled trial over a set of units indexed by $n]:=\{1,\ldots,n\}$. Each unit is assigned to either treatment ($T$) or control ($C$). The vector ${\bm{d}}\in{\mathbb{R}}^{n}$ denotes the treatment assignment, with $d_{i}=1$ if unit $i$ is in treatment and $d_{i}=0$ if it is in control. We also define ${\bm{z}}=2({\bm{d}}-1/2)$, so that $z_{i}=1$ for treated units and $z_{i}=-1$ otherwise. Finally, we define $\boldsymbol{q}$ as the vector with $i$-element equal to $q_{i}=p_{i}d_{i}+(1-p_{i})(1-d_{i})$.
For each unit $i$, the potential outcomes under treatment and control are denoted by $y^{(1)}_{i}$ and $y^{(0)}_{i}$, respectively. We treat these potential outcomes as fixed quantities; the only source of randomness in our setting is the treatment assignment. The observed outcome for unit $i$ is
@@
Each unit $i\in[n]$ has a fixed vector of pretreatment covariates ${\bm{x}}_{i}\in{\mathbb{R}}^{k}$. The matrix ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{n})^{\top}\in{\mathbb{R}}^{n\times k}$ collects the covariates for all units. For simplicity of exposition, we assume throughout the paper that $\bm{X}$} has full column rank. However, since our estimators are based on ridge regression, the results extend directly to cases in which $\bm{X}$} is rank-deficient.
We consider two standard treatment assignment mechanisms. Under simple random assignment, each unit is independently assigned to treatment with probability $p_{i}\in(0,1)$. Let $\boldsymbol{p}=(p_{1},\ldots,p_{n})^{\top}$ denote the vector of assignment probabilities, and define
$m=\min_{i\in[n]}\min\{p_{i},\,1-p_{i}\}.$
Under complete random assignment, the number of treated units is fixed at $n_{T}\in[n-1]$, and the number of control units is $n_{C}=n-n_{T}$. Treatment is assigned by selecting uniformly at random from all possible subsets of $n_{T}$ units in the sample.
Under simple random assignment, the Horvitz–Thompson (HT) estimator (horvitz1952generalization) is defined as

### v2

@@ Abstract
This article introduces a leave-one-out regression adjustment (LOORA) for estimating average treatment effects in randomized controlled trials. In finite samples, LOORA removes the bias of conventional regression adjustment and yields exact variance formulas for regression-adjusted Horvitz–Thompson and difference-in-means estimators. Ridge regularization curbs the influence of high-leverage observations, improving stability and precision in small samples. In large samples, LOORA matches the variance of the regression-adjusted estimator in [W. Lin (2013)] while remaining exactly unbiased. Two within-subject experimental applications, each providing a realistic joint distribution of potential outcomes as ground truth, show that LOORA removes substantial bias and achieves confidence interval coverage close to the nominal level.
@@ 1 Introduction
Establishing causal relationships is a central goal in economics and the social sciences. In observational studies, selection bias often obscures the effects of interventions. Randomized controlled trials (RCTs) address this problem by eliminating systematic pre-randomization differences between treatment and control groups. For this reason, RCTs occupy a central place in modern empirical research on causal inference.
@@
Most formal guarantees for regression-adjusted estimators rely on asymptotic arguments. However, early-stage clinical trials, marketing geo-experiments, and studies of firm- or country-level policies frequently operate in small sample regimes, where standard regression-adjusted estimators become biased, with undesirable consequences for treatment effect estimation and policy design (Freedman, 2008b). Moreover, in small samples, high-leverage observations (that is, observations whose covariate values exert a disproportionate influence on the regression fit) further degrade the performance of regression adjustment (Young, 2019).
@@
To address these challenges, we propose a regression adjustment framework that is unbiased in finite samples and robust to influential observations. A key component of our analysis is a leave-one-out regression adjustment (LOORA) procedure, which removes the bias of classical regression adjustment. Within this framework, we develop two estimators of the average treatment effect (ATE): LOORA-HT, a Horvitz–Thompson estimator for simple random assignment; and LOORA-DM, a difference-in-means estimator for complete random assignment. LOORA-HT and LOORA-DM are easy to implement and deliver a performance that previously required complex assignment mechanisms (see Harshaw et al., 2024).
@@
LOORA-HT and LOORA-DM simultaneously address three core problems in regression-adjusted estimation of treatment effects. First, they answer the critique of Freedman (2008a) by providing unbiased, regression-adjusted estimators for RCTs. Second, they attain the asymptotic variance of the estimator in Lin (2013), which is the efficient variance among all linearly adjusted estimators under simple and complete random assignment. Third, by stabilizing estimation in the presence of influential observations, they reduce the sensitivity to leverage documented in Young (2019).
1.1 Related Work
The practice of regression adjustment in randomized experiments dates at least to the use of analysis of covariance (ANCOVA) in classical experimental design, where a linear adjustment removes the effects of pretreatment covariates to improve precision. [^1] See, for example, discussions in Fisher (1971, Chapter 9). Despite its pervasiveness, the formal large-sample properties of regression adjustment after randomization have been repeatedly questioned. In two influential critiques, Freedman (2008b, a) argue that ordinary least squares (OLS) adjustments can introduce bias when the regression model is misspecified, and can be less efficient than the unadjusted difference-in-means estimator, even asymptotically.
In response, Lin (2013) shows that, under complete random assignment, an OLS regression of outcomes on treatment, covariates, and all treatment-by-covariate interactions delivers an estimator that is consistent for the average treatment effect and asymptotically efficient among the class of linearly adjusted estimators. While these results clarify the large sample properties of regression adjustment after randomization, two important gaps remain: (i) Lin (2013) does not provide finite-sample unbiasedness guarantees for regression adjustment, and (ii) the interacted specification in Lin (2013) can be unstable in moderate samples because it increases the number of regressors by adding interaction terms and potentially inflating leverage scores and sensitivity to influential observations.
This article proposes two estimators, LOORA-HT under simple random assignment and LOORA-DM under complete random assignment, that (i) are exactly unbiased for the finite-population average treatment effect, (ii) admit closed-form variance expressions in finite samples, and (iii) are asymptotically efficient among linearly adjusted estimators, attaining the variance bound of Lin (2013). We restrict our attention to linear regression adjustments because of their ubiquity in empirical research and the potential instability of nonlinear adjustments in small samples. Linear adjustments, however, can accommodate nonlinear relationships through covariate transformations.
LOORA combines leave-one-out regression adjustment and ridge regression to remove the finite-sample bias of regression-adjusted estimators and to curb the influence of high-leverage observations. For the difference-in-means setting, we show that LOORA-DM admits a representation that is equivalent to a leave-two-out construction in the sense of Spiess (2025), thereby satisfying necessary conditions for unbiasedness under complete random assignment. In this way, our estimators provide constructive finite-sample analogues of the asymptotic guarantees in Lin (2013), while remaining valid without the assumption of correctly-specified parametric outcome models.
@@ 1.1 Related Work
A related line of work examines precision gains in randomized studies through experimental design rather than regression adjustment. Harshaw et al. (2024) proposes a Gram–Schmidt walk design, which chooses treatment assignments in a way that approximately balances covariates. Their goal is to optimize the randomization design to deliver low-variance Horvitz–Thompson estimators in finite samples. In contrast, we consider simple or complete random assignment and instead adjust outcomes through leave-one-out regressions. This distinction matters in practice, because our methods can be applied to commonly used experimental designs without the need to change the assignment rule.
Ghadiri et al. (2023) introduces a cross-fitted, regression-adjusted Horvitz–Thompson estimator under simple random assignment. It splits the sample into two folds, estimates regression coefficients on one fold, and applies those coefficients to adjust outcomes in the other. This construction yields an exactly unbiased estimator with an explicit nonasymptotic variance bound. However, the bound is looser than the Lin (2013) asymptotic efficiency bound, and the article does not provide a procedure for constructing confidence intervals. Lei and Ding (2021) proposes a bias-correction procedure for regression adjustment when the number of covariates grows with the sample size. The method removes higher-order bias terms but does not yield exact unbiasedness in finite samples. Chang et al. (2024) develops an exact bias-correction approach that eliminates the bias, but does not derive the finite-sample variance of the resulting estimator. Gu et al. (2025) studies regression adjustment with high-dimensional covariates under a superpopulation framework and develops consistent estimators in this setting. Zhao et al. (2024) considers a related framework and proposes a debiasing approach based on higher-order influence functions that corrects own-observation bias arising from regression adjustment. Their theoretical development primarily focuses on unregularized linear adjustment, or regimes in which the ridge regularization parameter vanishes asymptotically, which is equivalent to using the Moore–Penrose pseudoinverse of the Gram matrix in the associated regression problem. In contrast, our analysis accommodates arbitrary regularization parameters. This flexibility enables explicit damping of high-leverage observations, which can otherwise induce unstable or unreliable inference, especially in finite-population or small-sample settings.
@@
A large concurrent literature studies regression adjustment and inference in more structured designs: stratified or blocked experiments, matched pairs, cluster randomization, or covariate-adaptive randomization. For example, Bai et al. (2024) survey modern design-based analysis of randomized experiments, emphasizing the role of stratification, regression adjustment and cluster randomized experiments. Bai et al. (2025b, a) develop design-based variance estimators and inference procedures for finely stratified and matched-pair experiments, including settings with imperfect compliance. Cytrynbaum (2024) characterizes asymptotically optimal linear covariate adjustment for general stratified randomization schemes.
Our approach provides an appealing alternative in settings where stratification becomes impractical or infeasible. In particular, stratification requires covariate information at the design stage, which is often unavailable in online experiments and field experiments with asynchronous enrollment. By contrast, simple and complete randomization designs apply naturally to these settings.
Section˜5 provides a more detailed comparison of our estimators with related methods.
@@ 1.2 Notation
We denote matrices and vectors by bold uppercase and lowercase letters, respectively. The $i^{\text{th}}$ entry of a vector $\bm{u}$} is denoted by $u_{i}$. The transposed $i^{\text{th}}$ row of a matrix $\bm{X}$} is denoted by ${\bm{x}}_{i}$, and its $(i,j)^{\text{th}}$ entry by $x_{ij}$. For a constant $c$ and a vector $\bm{u}$}, expressions such as $c+{\bm{u}}$ and ${\bm{u}}^{-1}$ are interpreted entrywise. When $\bm{u}$} and $\bm{v}$} are vectors of equal dimension, ${\bm{u}}{\bm{v}}$ and ${\bm{u}}/{\bm{v}}$ denote entrywise (Hadamard) product and division, respectively. A vector’s associated diagonal matrix appears in uppercase; for instance, $\boldsymbol{T}$ denotes the diagonal matrix $\bm{t}$} as its main diagonal. For any vector ${\bm{v}}\in{\mathbb{R}}^{n}$, the notation ${\bm{v}}_{-i}$ refers to the vector in ${\mathbb{R}}^{n-1}$ obtained by removing the $i^{\text{th}}$ entry of $\bm{v}$}. Likewise, ${\bm{X}}_{-i}$ denotes the matrix obtained by deleting the $i^{\text{th}}$ row of $\bm{X}$}. We denote by $\boldsymbol{1}$ the vector with all entries equal to one. For a matrix ${\bm{X}}\in{\mathbb{R}}^{n\times k}$, let $\overline{{\bm{X}}}$ denote the matrix whose rows are all equal to the average row of $\bm{X}$}, that is, each row of $\overline{{\bm{X}}}$ equals $\boldsymbol{1}^{\top}{\bm{X}}/n$. Similarly, for a vector ${\bm{y}}\in{\mathbb{R}}^{n}$, let $\overline{{\bm{y}}}$ denote the vector whose entries are all equal to the average of the entries in $\bm{y}$}.
Norms and asymptotic notation. For a vector ${\bm{u}}\in{\mathbb{R}}^{n}$, $\left\|{\bm{u}}\right\|_{2}=\sqrt{{\bm{u}}^{\top}{\bm{u}}}$ denotes the Euclidean norm, and $\left\|{\bm{u}}\right\|_{\infty}=\max_{i\in[n]}|u_{i}|$ denotes the $\ell_{\infty}$ norm. The operator ${\rm tr}(\cdot)$ denotes the trace of a square matrix. For a matrix ${\bm{X}}\in{\mathbb{R}}^{n\times k}$, $\left\|{\bm{X}}\right\|_{F}=\sqrt{{\rm tr}({\bm{X}}^{\top}{\bm{X}})}$ denotes the Frobenius norm, and $\left\|{\bm{X}}\right\|_{2,\infty}=\max_{i\in[n]}\left\|{\bm{x}}_{i}\right\|_{2}$ denotes its $(2,\infty)$-operator norm. Let $\mathbb{N}$ be the set of natural numbers. For any two functions $f,g\colon\mathbb{N}\to{\mathbb{R}}$, we write $f(n)=O(g(n))$ if there exist constants $C>0$ and $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$, $\lvert f(n)\rvert\leq C\lvert g(n)\rvert$. Similarly, we write $f(n)=\Omega(g(n))$ if there exist constants $C>0$ and $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$, $\lvert f(n)\rvert\geq C\lvert g(n)\rvert$.
Projection matrices. Consider a full-rank matrix ${\bm{X}}\in{\mathbb{R}}^{n\times k}$ with $n\geq k$. We denote the projection matrix by $\boldsymbol{H}={\bm{X}}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{X}}^{\top}$, and its diagonal entries by $h_{ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{x}}_{i}$, which represent the leverage scores of rows $i\in[n]$. The identity matrix is denoted by $\bm{I}$}. For $\lambda\geq 0$, the ridge projection matrix is $\boldsymbol{H}_{\lambda}={\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}$, whose diagonal entries $h_{\lambda ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}$ are the ridge leverage scores.
@@ 2 Finite Population Framework
We operate within the Neyman–Rubin potential outcomes framework (Neyman, 1923; Rubin, 1974) for a randomized controlled trial over a set of units indexed by $n]=\{1,\ldots,n\}$. Each unit is assigned to either treatment or control. The vector ${\bm{d}}\in{\mathbb{R}}^{n}$ denotes the treatment assignment, with $d_{i}=1$ if unit $i$ is in treatment and $d_{i}=0$ if it is in control. We also define ${\bm{z}}=2({\bm{d}}-1/2)$, so that $z_{i}=1$ for treated units and $z_{i}=-1$ otherwise. Finally, we define $\boldsymbol{q}$ as the vector with $i^{\text{th}}$ element equal to $q_{i}=p_{i}d_{i}+(1-p_{i})(1-d_{i})$, where $p_{i}$ is the treatment-assignment probability for unit $i$.
For each unit $i$, the potential outcomes under treatment and control are denoted by $y^{(1)}_{i}$ and $y^{(0)}_{i}$, respectively. We treat potential outcomes as fixed quantities; the only source of randomness in our setting is the treatment assignment. The observed outcome for unit $i$ is
@@
Each unit $i\in[n]$ comes with a vector of fixed pretreatment covariates ${\bm{x}}_{i}\in{\mathbb{R}}^{k}$. The matrix ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{n})^{\top}\in{\mathbb{R}}^{n\times k}$ collects covariate values for all units. For simplicity of exposition, we assume throughout the article that $\bm{X}$} has full column rank. However, since our estimators are based on ridge regression, the results extend directly to cases in which $\bm{X}$} is rank-deficient.
We consider two standard treatment assignment mechanisms. Under simple random assignment, each unit is independently assigned to treatment with probability $p_{i}\in(0,1)$, with $\boldsymbol{p}=(p_{1},\ldots,p_{n})^{\top}$. Under complete random assignment, the number of treated units is fixed at $n_{T}\in[n-1]$, and the number of control units is $n_{C}=n-n_{T}$. Treatment is assigned by selecting uniformly at random from all possible subsets of $n_{T}$ units in the sample.
Under simple random assignment, the Horvitz–Thompson (HT) estimator (Horvitz and Thompson, 1952) is defined as